OpenStudy (anonymous):

find the derivative of 2(x^2 + y^2)^2 = 25(x^2 - y^2)

OpenStudy (anonymous):

Make implicit derivative

OpenStudy (anonymous):

that is: 2*2*(x^2+y^2)*(2x+2y*y') = 25*(2x-2y*y')

OpenStudy (anonymous):

now just isolate y'

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

y', isolate y' now

OpenStudy (anonymous):

i kknow but there are 5 ys

OpenStudy (anonymous):

im just kidding

OpenStudy (zale101):

Treat y' as another variable and you're trying to solve for it, doing simple algebra basically.

OpenStudy (anonymous):

to isolate y, you must first solve the parentheses

OpenStudy (zale101):

@cleetus779 do you have any idea on what you're supposed to do?

OpenStudy (anonymous):

yes, working it out, and came up with a slightly different answer than wolfram, algebraic mistake

OpenStudy (anonymous):

got this after multiplying: 8x^3 + 8y^2x + 2yx^2(dy/dx) + 8y^3(dy/dx) = 50x + 50y(dy/dx)